Two planets are 3.4 x 10^10 m apart (centre to centre) in a straight line far from the influence of any other objects. The first planet has a mass of 2.5 x 10^22 kg. The second planet has a mass of 8.9 x 10^21 kg. How far from the centre of the first planet is the overall acceleration due to gravity, g, equal to zero?
Answer #1

let us consider that point is at distance d from 1st planet

also, at that point acceleration due to gravity due to each particle must be same

so

G M1 / d^2 = G M2 / ( 3.4 - d) ^2

3.4 / d - 1 = sqrt ( 8.9 / 25)

d = 2.129* 10^10 m

=======

Comment before rate in case any doubt, will reply for sure.. goodluck